PFI-ZEKE PHOTOELECTRON SPECTROSCOPY OF METASTABLE He_{2}: IONIZATION POTENTIAL AND ROVIBRATIONAL STRUCTURE OF He_{2}^{+}

\underline{MARTIN SCHÄFER}, MATTHIAS RAUNHARDT, NICOLAS VANHAECKE^{a}, and FRÉDÉRIC MERKT, ETH Zürich, Laboratorium für Physikalische Chemie, Wolfgang-Pauli-Strasse 10, 8093 Zürich, Switzerland.

A supersonic beam of metastable He^{∗} atoms and He_{2}^{∗} a ^{3}Σ_{u}^{+} molecules has been generated using a pulsed discharge at the exit of a pulsed valve prior to the gas expansion into vacuum. Pulsed-field-ionization zero-kinetic-energy (PFI-ZEKE) photoelectron spectra of the He_{2}^{+} X^{+} ^{2}Σ_{u}^{+} (v^{+} =0 − 2) ← He_{2}^{∗} a ^{3}Σ_{u}^{+} (v′′ =0 − 2) transitions and photoionization spectra of He_{2}^{∗} in the vicinity of the lowest ionization thresholds have been recorded. The energy level structures of ^{4}He_{2}^{+} X^{+} 2Σ_{u}^{+} (v^{+} ≤ 2, N^{+} ≤ 23) and ^{3}He_{2}^{+} X^{+} ^{2}Σ_{u}^{+} (v^{+} = 0, N^{+} ≤ 11) have been determined, and an accurate set of molecular constants for all isotopomers of He_{2}^{+} has been derived in a global analysis of all spectroscopic data reported to date on the low vibrational levels of He_{2}^{+}.^{b} The analysis of the photoionization spectrum by multichannel quantum defect theory has provided a set of parameters describing the threshold photoionization dynamics of He_{2}^{∗} a ^{3}Σ_{u}^{+}.

aPermanent address: Laboratoire Aimé Cotton, CNRS II, Bâtiment 505, Campus d’Orsay, 91405 Orsay cedex, France

bD. S. Ginter, M. L. Ginter, J. Mol. Spectrosc. 82, 152–175 (1980). N. Yu and W. H. Wing, Phys. Rev. Lett. 59, 2055–2058 (1987), Phys. Rev. Lett. 60, 2445 (1988).